OpenStudy (anonymous):

how do I set up the integral to find the volume of the solid obtained by revolving the region about the x-axis when y=square root of x, y=1 and x=4

OpenStudy (anonymous):

I think it turns into a disk but I am not sure. And I cant figure out the integral.

OpenStudy (dumbcow):

 this uses the washer method where for each circular cross-section you take large area and subtract smaller area Here the outer radius is sqrt(x), the inner radius is 1 \[V = \pi \int\limits_{?}^{?}R^{2} -r^{2} \] \[V = \pi \int\limits_{1}^{4}(\sqrt{x}^{2} - 1^{2}) dx = \pi \int\limits_{1}^{4}(x - 1)dx\]